Assume that in country A, the unit labor requirement for producing good X is 100 hours, and the unit labor requirement for good Y is 20 hours. Meanwhile in country B, the unit labor requirement for producing good X is 80 hours and the unit labor requirement for good Y is 40 hours. Answer the following:

  1. Explain which country has an absolute advantage in each good.
  2. Explain which country has a comparative advantage in each good.
  3. Under the Theory of Comparative Advantage, explain the trade pattern between these countries.
  4. Draw a Production possibilities curve for both countries (2 graphs, one for each country), given the above data. Show and explain how international trade could leave both countries better off than without any trade. Explain.
Answer #1

Unit labor requirement table

In hours

X Y
A 100 20
B 80 40

1) absolute advantage is related to productivity

So B takes only 80 hours to produce one X, as against 100 hours in case of nation A

Thus B has absolute advantage in X

& A has absolute advantage in Y

2) comparative advantage

Opportunity cost of production of X in A = 100/20 = 5 units of Y

In B, OC of production of X in B = 80/40 = 2 units of Y

So as B has lower OC of production of X in terms of Y

So B has Comparative advantage in X

& A has comparative advantage in Y

.

3) trade pattern

Nation will export the good in which it has comparative advantage

So ​​​​B will export X

& A will export Y

.

4)

PPC , downward sloping

Good X on X axis

Slope of nation A PPC= opportunity cost of X = 5

In B

PPC slope = 2

​​​​​